OpenStudy (anonymous):

[9.01] Identify the vertex for the graph of y = −3x2 − 6x + 4. (1, 13) (−1, 13) (1, 7) (−1, 7)

OpenStudy (anonymous):

there are two ways I can think to solve this, each depend on what math class you are in, algebra or calc?

OpenStudy (anonymous):

algebra

OpenStudy (jdoe0001):

for a parabola equation like \(\bf ax^2+bx+c\) you can find the the vertex at the coordinates of \(\bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=-3x^2-6x+4 y=-3(x^2+2x)+4 Factor -3 y=-3(x^2+2x+1)+4+3 Complete square y=-3(x+1)^2+7 vetex (-1,7)

OpenStudy (anonymous):

Oh ok so its D

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you can you help me on this [9.01] What is the value of the coefficient "b" when the quadratic equation y = (3x + 2)(2x − 1) is written in standard form? −7 −2 1 6

OpenStudy (anonymous):

will it be -7

OpenStudy (anonymous):

its not -7, foil the equation and you'll see what 'b' becomes

OpenStudy (anonymous):

No its not -7 its 1 http://openstudy.com/study#/updates/50719bbfe4b0c2dc8340dcb1 Where I found the answer

OpenStudy (anonymous):

y = (3x+2)(2x-1) y = 6x^2 +4x -3x - 2 y = 6x^2 +x -2 y = ax^2 +bx + c

OpenStudy (anonymous):

It was easy I just over thought it